A sample containing a mixture of tin (ii)oxide and tin(iv)oxide and weighing 2.150g is dissolved in dilute acid.The tin(ii) is converted to Sn⁴+ and tin(iv) converted to Sn²+.This solution is titrated with 0.1500M K2C2O7 potassium dichromate (vi) solution and 18.55cm³ of dichromate solution are required to reach the end point,hw many moles of SnO was titrated.
1
Expert's answer
2019-02-07T05:29:38-0500

C1V1 = C2V2

n(K2Cr2O7) = C*V = 0.15mol/L*0.01855L = 0.00278 mol

n(SnO) = 0.00278 mol